Please only answer if 100% sure I keep getting incorrect responsesSuppose that an economy is characterized by M-$10 trillion V- 1.8 P-base index 1.0 nstructions: Enter your responses rounded

Suppose that an economy is characterized by M-$10 trillion V- 1.8 P-base index 1.0 

Instructions: Enter your responses rounded to two decimal places (do not include a negative sign ()with your answers). 

a. What is the real value of output (Q)? $18 trillion Now assume that the Fed increases the money supply by 20 percent and velocity remains unchanged 

b. If the price level remains constant, by how much will real output increase? 

c. If, instead, real output is fixed at the natural level of unemployment (-Q from part a), by how much will prices increase in percentage terms? 

d. By how much would Vhave to decrease to offset the increase in M (assuming Qand P did not change)?
Answer #1


(a)

The equation of exchange is as follows -

M * V = P * Q

$10 trillion * 1.8 = 1 * Q

Q = $18 trillion

Thus,

The real value of output (Q) is $18 trillion.

(b)

Now, money supply increases by 20 percent.

New money supply = $10 trillion + ($10 trillion * 0.20) = $10 trillion + $2 trillion = $12 trillion

Price level and velocity remains unchanged.

So,

M * V = P * Q

$12 trillion * 1.8 = 1 * Q

$21.6 trillion = Q

Thus,

The real output increases by $3.6 trillion.

(c)

New money supply = $12 trillion

Velocity and real output remains unchanged.

So,

M * V = P * Q

$12 trillion * 1.8 = P * $18 trillion

$21.6 trillion = P * $18 trillion

P = $21.6 trillion/$18 trillion

P = 1.2

Increase in price level = [(1.2 - 1)/1] * 100 = 20%

The price level will increase by 20%.

(d)

M = $12 trillion

P = 1

Q = 18 trillion

M * V = P * Q

$12 trillion * V = 1 * $18 trillion

V = $18 trillion/$12 trillion

V = 1.5

Thus,

The V would have decreased by 0.3